OpenStudy (anonymous):

4 1/6 - 2 5/6 Subtract. Write the answer as a mixed number in simplest form.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

like a guess?

OpenStudy (anonymous):

8/3 but i dnt know about the mixed part someone else prob can get it right

OpenStudy (johnweldon1993):

Alright we can change mixed numbers to improper fractions and then change back... We have right now \[\large 4\frac{1}{6} - 2\frac{5}{6}\] To change those mixed numbers to improper fractions we take the denominator, multiply it by the whole number...and then add the denominator to that result. Then we just place that number over the original denominator So \[\large 4\frac{1}{6} = \frac{4 \times 6 + 1}{6} = \frac{25}{6}\] and \[\large 2\frac{5}{6} = \frac{6 \times 2 + 5}{6} = \frac{17}{6}\] So we have \[\large \frac{25}{6} - \frac{17}{6} = \frac{8}{6}\]

OpenStudy (johnweldon1993):

To change that BACK to a mixed number, we simply do the division \[\large \frac{8}{6}\] how many times does 6 go into 8? just once (this will be your whole number) How much is left over from that? well since 6 only goes in once, that is 6 out of the 8...so there is a remainder of 2 (this will be our numerator) and finally the denominator still stays the same so... \[\large \frac{8}{6} = 1\frac{2}{6} = 1\frac{1}{3}\]